A body travels 2m in the 2nd second and 6m in the next 4 seconds. What will be the distance travelled in 9th second.

Grade:11

3 Answers

Vasantha Kumari
10 years ago

Solution:

As per the given data we have distance covered in 2nd sec = 2m and distance travelled is 6m in 4 sec.

So, at the end of 9th sec the distance would be 9m.

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty.

Nirmal Singh.
10 years ago

Let's consider initial velocity is u and acceleration is a and 8th sec velocity is v..

we will this equation S=ut+0.5at2 as shown below

2 = u*2 +0.5 *a*2*2 -----1
8 = u*6 + 0.5*a*6*6 ------2

now 3x(2 = u*2 +0.5 *a*2*2 ) ----3

now we will subtract from equation 3 to equation 2

we will get a = 1/6 m/s^2

Putting above value of a in eq 2 we get

8 = u*6+0.5*1/6*6*6

u = 5/6 m/s^2

v = u+at


Now, v = 5/6+ 1/6*9
So, v=7/3 m/s

Thanks & Regards,

Nirmal Singh

Askiitians Faculty

Prashant
5 years ago
1 ) 2s it covers 2m .
2) 6s (next 4 sec means 2+4) it covers 8m (2m+6m) ...
compare the ratios ...
6s/2s : 8m/2s we get, 3:4 ...
now to find dist travelled in 9s x/4=9/3 x=12m